### v1

# The unreasonable likelihood of being Origin of life, terraforming, and AI

Abstract
The origin of life on Earth via the spontaneous emergence of a protocell prior to Darwinian evolution remains a fundamental open question in physics and chemistry. Here, we develop a conceptual framework based on information theory and algorithmic complexity. Using estimates grounded in modern computational models, we evaluate the difficulty of assembling structured biological information under plausible prebiotic conditions. Our results highlight the formidable entropic and informational barriers to forming a viable protocell within the available window of Earth’s early history. While the idea of Earth being terraformed by advanced extraterrestrials might violate Occam’s razor from within mainstream science, directed panspermia—originally proposed by Francis Crick and Leslie Orgel—remains a speculative but logically open alternative. Ultimately, uncovering physical principles for life’s spontaneous emergence remains a grand challenge for biological physics.
Beginnings
“All cells come from cells” [1] leads us into a classic chicken-and-egg dilemma: where did the first cell come from? Either it came from somewhere else—conveniently outsourcing the mystery—or it emerged from the laws of physics and chemistry on a young, chaotic, and geologically active Earth. Once a minimal protocell or replicator emerged, Darwinian evolution could take over, merely selecting and diversifying its way toward greater complexity, biodiversity, and eventually, the modern biosphere, from gold fish to particle accelerators and social media. One thing this onset of a discussion illustrates - how little we know about our origin, and the cosmos in general.
As for the “somewhere else” hypothesis, science fiction has long entertained terraforming scenarios. In 2001: A Space Odyssey, an inscrutable black monolith gives early hominins a mysterious push toward cognition—and tool use [2]. In Star Trek, entire planetary ecosystems are engineered by advanced civilizations or errant AI, often with remarkable speed [3]. The X-Files proposed that alien viruses might have seeded Earth, albeit oddly fixated on rural America [4]. In Prometheus, a lone extraterrestrial—credited only as “the Engineer”—dissolves himself in a waterfall, presumably catalyzing Earth’s biochemistry, though the narrative coherence dissolves nearly as fast [5]. The mother of all terraforming stories is, however, the Book of Genesis, written more than 2,500 years ago. It declares that life was created in six days [6] — a claim that has profoundly shaped human thought, even if not amenable to scientific testing.
One step closer to science (though not necessarily to certainty), Francis Crick—the co-discoverer of DNA’s structure—and Leslie Orgel proposed in 1973 the idea of directed panspermia [7]. In their scenario, an advanced extraterrestrial civilization, facing extinction or perhaps scientific curiosity, dispatches microbial “starter kits” to habitable planets like ours (see Fig. 1 for an illustration). While Crick and Orgel attempted to formulate this idea more like a testable hypothesis, it deftly relocates the explanatory burden to someone else’s biochemistry.
Meanwhile, modern Bayesian analyses of life’s emergence struggle with deeply unconstrained data (or lack thereof): a sample size of one, and priors shaped not only by empirical observation but also by immense subjectivity [8, 9]. There’s also the problem of timing—life appeared early on Earth, yet the intelligent observer required to marvel this fact arrived very much later, introducing a strong observational bias. In this work, we revisit the origin of life through the lens of modern cell models and AI-based tools, to better appreciate the informational and physical complexity required for protocell formation. We begin with general considerations and propose a qualitative theoretical framework. Finally, we follow Feynman’s mantra that ”doubt and uncertainty are not to be feared but welcomed” [10], and attempt to cautiously conclude something.
Figure 1: Fantasy sci-fi imagery of terraforming. Chatgpt4.0’s hallucination of early Earth with seeded biomaterial, jump starting Darwinian evolution.
Geological constraints can be soft
When did life emerge, and what are the soft and hard constraints? There were two massive impacts with global sterilizing effects: Earth was first struck by a Mars-sized body, named “Theia,” about 4.51 Gy ago, leading to the formation of the Moon [11]. This was followed by another sterilizing impact, dubbed “Moneta,” approximately 40 My later [12]. Mineralogical evidence from zircons indicates that both an atmosphere and liquid water were already present on Earth’s surface around 4.404 Gy ago [13]. The oldest putative evidence for life comes from 13C-depleted carbon inclusions in 4.1 Gy-old zircon deposits [14]. The oldest direct and undisputed evidence is found in microfossils preserved in 3.465 Gy-old rocks in Western Australia [15, 16, 17].
Recently, a study combining molecular clock methodology, horizontal gene-transfer-aware phylogenetic reconciliation, and existing biogeochemical models addressed the nature of the last universal common ancestor (LUCA), estimating its age, gene content, metabolism, and ecological context [18]. Surprisingly, LUCA is estimated to have lived around $\sim 4.2$ Gy ago, based on calibration with fossil and isotope records. More perplexing still, LUCA appears to have been an anaerobic acetogen, metabolically similar to modern prokaryotes, and already equipped with ATP synthesis, a TCA cycle, early immune capabilities including CRISPR-Cas effector proteins, and embedded within a microbial ecosystem. A recent consensus across eight genomic and proteomic studies broadly supports this view [19]. This suggests that a protocell—presumably far simpler—must have emerged even earlier, perhaps relying on RNA or other primitive replicator systems.
Only in the last blink of planetary time — some 100,000 years ago — did anatomically modern humans appear, wielding curiosity and stunning technological inventions. Although the controlled use of fire by earlier hominins dates back about 1 Myr [20], in geological terms intelligence has arrived extraordinarily late. Yet anthropic considerations must be taken into account: without human observers, it would not have been possible to ponder this spectacle. Intriguingly, this late arrival may be further constrained from the other end of the timeline, as Earth’s biosphere is projected to collapse in roughly 100 Myr due to impending “CO2 starvation” (even way before the sun becomes a Red Giant and engulfs the Earth) [21]. As solar luminosity increases, silicate weathering accelerates, drawing down atmospheric CO2. Below $\sim 150$ ppm, photosynthesis ceases, triggering ecological collapse. Hence, if intelligent life requires too much time to emerge, many planets—especially around hotter, short-lived stars—may simply run out of time, biasing observations toward an unusually early start.
On a more optimistic note, a modified Drake equation—updated with exoplanet data—places a firm lower bound on the likelihood that at least one other technological species has arisen at some point in the observable universe. If the probability that a habitable-zone planet spawns a technological species is $\geq 10^{-24}$ — about the odds of winning the lottery a modest $8$ times in a row — then humanity is almost certainly not alone [22]. So statistically speaking, we’re probably not the only ones wondering whether we’re alone.
Physical principles of life’s emergence - anyone?
Since Erwin Schrödinger’s landmark 1944 essay What is Life?, physicists have speculated whether the emergence of life might require physical principles beyond those already known [23]. Schrödinger speculated—presciently, though vaguely—about undiscovered laws governing order and organization in living systems, prompting generations of researchers to explore the physics of self-organization. Since then, the field has progressed dramatically. Here is a short list without any claim of completeness:
Alan Turing’s work on reaction–diffusion systems introduced a mathematical framework for how spatial patterns, such as those observed in morphogenesis, can spontaneously arise from simple rules [24]. The Landauer–Bennett principle revealed a deep connection between information, entropy, and the thermodynamic cost of computation—key concepts for understanding living systems as information-processing entities [25, 26]. Ilya Prigogine’s theory of dissipative structures demonstrated how local order can emerge in open systems that export entropy [27]. Stuart Kauffman’s concept of autocatalytic sets offered a compelling model for how chemical networks might self-amplify and sustain themselves [28]. The latter two demonstrate nicely that complex systems have new emergent collective features beyond the sum of their parts, not predictable from first principles [29].
More recently, in work bordering the popular science literature, Jeremy England has proposed that life-like order could emerge generically through driven dissipation under physical selection-like conditions [30], while Sara Walker has emphasized the centrality of information flow and causal structure in defining what it means to be alive [31]. Both argue that to understand life’s origin, we must look beyond molecular inventories and instead focus on the dynamics of information, energy, and organization in physical systems. If life emerged through a relay of quasi-stable, far-from-equilibrium states — each more improbable than the last yet somehow energetically favored — then the transition becomes not miraculous, but thermodynamically admissible. The origin of life would simply be extension of the classical rule that ’exergonic reactions occur spontaneously’ [32].
Broadly speaking, origin-of-life relevant research falls within the rich and rapidly evolving domains of active matter, stochastic thermodynamics, and nonequilibrium phase transitions. These fields have significantly expanded our understanding of how systems far from equilibrium can maintain structure and perform work. Yet despite this progress, a fully satisfying mechanism for the spontaneous emergence of a protocell from disordered chemistry remains elusive.
Dogmas and artificial intelligence
In science, dogmas define domains of accepted knowledge, enabling structured peer review and funding decisions. Yet they can also inhibit new discoveries [33]. For example, the phrase “nothing in biology makes sense except in the light of evolution” [34], though foundational in lab work and bioinformatics, may inadvertently create boundaries that hinder unconventional thinking about life’s origins. Similarly, the dictum “all cells come from cells” [1] tends to obscure the deeper question of how the first cell came to be.
One thing is certain: with the rise of powerful AI tools and mechanistic models, we now have entirely new ways of exploring biological complexity. Tools like AlphaFold for protein folding [35] and comprehensive whole-cell models [36, 37, 38] allow us to estimate the information content of life using Kolmogorov (algorithmic) complexity. This measures the shortest description required to reproduce a system—the more structured or regular a system, the more it can be compressed and the easier it can be made [39]. Here, by combining such complexity estimates with basic rate-distortion theory [40, 41], we aim to shed new light on the plausibility of life’s rapid emergence.
Theoretical framework
We approach the problem of life’s origin by estimating how plausible it is for chemical processes to spontaneously assemble a sufficiently complex protocell within the available time on early Earth. We assume this window to be about $500$ Myr—interpolated between “optimistic” ($304$ Myr) and “conservative” ($939$ Myr) estimates [9]. We apply a heuristic version of Shannon’s rate-distortion theory [40, 41], which quantifies the minimal information rate $R(D)$ needed to construct a functioning system, given a tolerated distortion level $D$. Here, $D$ characterizes how much loss (error or degradation) can be tolerated in forming a viable protocell. The distortion rate depends on the entropy of the chaotic prebiotic environment, denoted $H_{\text{prebiotic}}$, which represents the information content of the chemical soup in a pond or hydrothermal vent.
To succeed, the rate $R(D)$ must be at least as large as the macroscopic information rate required to assemble a protocell of complexity $I_{\text{protocell}}$ within the time available, $T_{\text{available}}=500$ Myr (see Fig. 2). We begin by writing in mathematical form:
$R(D)\approx\underbrace{\eta\frac{H_{\text{prebiotic}}}{D}}_{\text{entropy collapse}}\geq\underbrace{R_{\text{min}}=\frac{I_{\text{protocell}}}{T_{\text{available}}}}_{\text{rise in complexity}}$ (1)
In this simplified formulation, we approximate $R(D)$, typically convex and decreasing, by the ratio of prebiotic entropy, $H_{\text{prebiotic}}$, and molecular persistence time, $D$, where $\eta$ is an efficiency prefactor (also called a ”fudge factor”). This term accounts for the fact that not all molecular encounters result in usable contributions. Molecules may be structurally misaligned, chemically unstable, or fail to integrate into a coherent assembly due to the absence of repair mechanisms or spatial constraints in the local environment.
Admittedly, all quantities in Eq. 1—except perhaps $T_{\text{available}}$—are highly uncertain. Nevertheless, following this line of reasoning, the spontaneous origin of life remains plausible if the left-hand side exceeds the right-hand side. If not, we may be missing some crucial ingredient—such as additional evolutionary pathways, unknown physical principles, or external interventions like Crick and Orgel’s directed panspermia (or, more colorfully, aliens terraforming the planet). For the sake of argument, we set aside the ”cheap” statistical fallback that rare events inevitably occur somewhere in a vast universe.
Figure 2: Mathematical framework based on rate-distortion theory. (A) Funnel of collapsing entropy for the emergence of a protocell. (B) Rate-distortion function from information theory, comparing $R(D)\sim H_{\text{prebiotic}}/D$ as a function of $D$ with $R_{\text{min}}$ from observational constraints (see the main text for explanations).
Entropy of prebiotic soup and tolerated error rate of assembly
In the absence of a time machine, we estimate conditions on Earth roughly 4 Gy ago. We begin with the left-hand side of Eq. 1, aiming to estimate the entropy of the chemical soup, $H_{\text{prebiotic}}$, and the tolerated error rate $D$—which characterizes how long complex organic molecules persist before degrading. Importantly, the entropy $H_{\text{prebiotic}}$ pertains not to an entire ocean or vent system, but to the volume accessible to a nascent protocell. While the volume of a single protocell may be around 1 femtolitre ($10^{-15}$ l), relevant molecules could diffuse in from a larger surrounding droplet. We seek order-of-magnitude estimates informed by chemical experiments simulating early Earth conditions, meteorite data, and planetary spectroscopy.
In the Miller–Urey experiments, spark discharges in a reducing gas mixture produced a variety of amino acids and other small organic molecules [42]. During the Cassini–Huygens mission, analysis of Titan’s atmosphere revealed that complex photochemistry can generate thousands of distinct organic species, including nitriles, hydrocarbons, and possible prebiotic precursors [43]. Similarly, carbonaceous chondrite meteorites (e.g., Murchison) contain more than 60 amino acids, as well as nucleobases, polyols, carboxylic acids, and related compounds [44]. This is consistent with samples returned from asteroid Bennu by the OSIRIS-REx mission, which revealed a rich inventory of organic compounds—including those preserved in salt-rich brine inclusions in phyllosilicate minerals, dating back to $\sim 4.5$ billion years [45].
Combinatorial chemistry across diverse geochemical settings could plausibly yield thousands of unique organic molecules, suggesting a chemical library on the order of $10^{5}$$10^{6}$ small molecules in the prebiotic environment. Assuming an average information content of $\sim 10$ bits per molecule—representing functional or structural specificity—this allows for roughly $10^{3}$ distinguishable states per molecule (e.g., different side chains, folding propensities, or reactivities). This estimate aligns with protein domain annotations and bits-per-residue values from sequence-function maps [46, 47, 48], as well as compression-based descriptors like SMARTS or InChI [49]. Multiplying, we estimate the prebiotic entropy as $H_{\text{prebiotic}}\sim 10^{6}$$10^{7}$ bits.
Next, we consider the distortion timescale $D$, interpreted as the average lifetime of a complex organic molecule before degradation by hydrolysis, UV damage, oxidation, or further transformation. This defines the “error rate” or persistence window for useful molecular structures. Under plausible early-Earth conditions, complex organics such as ribose and peptides have half-lives ranging from hours to days. For example, ribose degrades with a half-life of $\sim 3$ hours at pH 7 and 37C [50]. Small peptides degrade over hours to days depending on sequence and pH. Nucleobases like adenine are sensitive to solar UV and degrade within days to weeks. Without an ozone layer, degradation could be faster—on the scale of hours to days. Conservatively, we adopt a mean lifetime of $\sim 1$ day, or about $10^{5}$ seconds, for moderately stable molecules in a warm, aqueous, anoxic environment exposed to UV. However, in cryogenic or mineral-protected contexts—such as asteroid interiors—organic molecules may persist for billions of years [45]. The existence of such intergalatic freezers implies that $D$ could vary over many orders of magnitude depending on local geochemical context.
Combining $H_{\text{prebiotic}}\sim 10^{6}$$10^{7}$ bits with $D\sim 10^{5}$ s and assuming optimistically an efficiency $\eta=1$, we obtain:
[…]
Thus, chemical evolution must extract or preserve about 100 bits per second to successfully cross the informational threshold required for life’s emergence.
For further intuition, here are two aids helping to think about the meaning of this estimate. Consider, if you like, a pool of $10^{7}$ bits of potentially useful molecules, each lasting about $10^{5}$ seconds. To prevent this information from vanishing forever, you need to stabilize or incorporate about 100 bits per second. Or more dramatically, picture a ‘melting library’: you stand in a library with 10 million books ($H_{\text{prebiotic}}$), each self-destructing 24 hours after removal. Your task is to compile a manual for “building life.” To capture information before it disappears, you’d need to scan roughly 100 books per second.
If the molecular lifetime were much shorter—say $D\sim 1$ s, due to rapid diffusion or lack of protective mechanisms—then the required information rate would increase accordingly. The key idea remains: if $R(D)$ is too low to meet $R_{\text{min}}=I_{\text{protocell}}/T_{\text{available}}$, then abiotic emergence becomes implausible without invoking directed evolution, external contributions (e.g., panspermia), or unknown self-organizing principles. The delivery of relatively stable, complex molecules—by meteorites or interstellar dust—could act as a bootstrap for prebiotic evolution [45]. Given the large uncertainties in parameter values, our aim is not to deliver a definitive conclusion but to structure a quantitative discussion.
Information content of a protocell
The right-hand side of Eq. 1 requires an estimate of the complexity of a protocell, $I_{\text{protocell}}$. How can we quantify this complexity, including its structural and dynamic organization? It shouldn’t come as a surprise that cells are not simply bags of well-mixed molecules, but highly structured systems—down to the folding and interaction of their constituent proteins. Moreover, cellular processes are orchestrated in time. If proteins are the musicians, then signaling, metabolic, and genetic pathways are the conductor and the score — guiding timing, harmony, and transitions.
Nature’s “code” is embedded in physical law and can often be expressed in elegant mathematical equations—sometimes compact enough to fit on a single sheet of paper. Unlike our computational models, Nature does not compute itself; it simply unfolds. To make sense of this complexity, we turn to symbolic and algorithmic representations, such as equations or computer code. Yet this translation from physical causality to epistemic representation introduces ambiguity: we cannot neatly separate what is essential from what is merely descriptive. As a result, estimates of Kolmogorov complexity are inherently fuzzy. From below, complexity is bounded by simple, universal laws; from above, it is constrained by the size of compressed code needed to realistically simulate the system. But even this upper bound is imperfect—it includes implementation overhead and excludes unknown biological details. In short, the more we compress reality, the more we risk distorting its true complexity.
To proceed, we divide the total information content of a protocell, somewhat artificially, into distinct components:
$I_{\text{protocell}}=I_{g}+I_{s}+I_{d}$
where $I_{g}$ is the genetic information, $I_{s}$ the structural information (e.g. protein folds, cellular organization), and $I_{d}$ the dynamic information encoded in biochemical pathways. Each term is estimated via Kolmogorov (algorithmic) complexity, defined as the length of the shortest program that can realistically generate the relevant component, at least conceptually.
The genomic information is the easiest to estimate—and, perhaps surprisingly, contributes the least. Consider a minimal self-replicating organism such as Mycoplasma genitalium, which contains roughly 500 genes [37], each about 1,000 base pairs long. With two bits per base pair (for A, T, C, G), this yields $I_{g}\sim 500\times 1000\times 2=10^{6}$ bits. This estimate is supported by the synthetic minimal cell JCVI-syn3.0, developed from Mycoplasma mycoides, which contains only 473 genes [51]. Remarkably, 149 of these genes still have unknown function—underscoring how much we have yet to understand, even in minimal life.
To estimate $I_{s}$, we consider protein structure and folding. A protein’s biological function depends on its shape. While AlphaFold2—an AI-based protein-folding algorithm—occupies 556 GB, this includes thousands of structures, redundancies, and metadata [35]. A more realistic lower bound for a minimal cell might be in the range of $10^{6}$$10^{8}$ bits, depending on the number of unique folds, protein domains, and the information needed to describe each.
To estimate $I_{d}$, the dynamic component, we turn to whole-cell models. The first comprehensive whole-cell simulation of Mycoplasma genitalium integrated 28 submodels covering DNA replication, transcription, translation, metabolism, and more [37]. The compressed code size was about 140 MB, or roughly $10^{9}$ bits. The model used over 1,900 parameters curated from 900+ publications and could predict phenotypes from genotypes. While this is almost certainly an overestimate—due to redundant code and parameters—it’s informative. Notably, the model excludes spatial organization, mechanical feedback, and memory effects. Other whole-cell simulations of E. coli [38] and synthetic cell JCVI-syn3A [52] use somewhat smaller codebases but follow similar principles.
Taken together, we estimate the total information content of a protocell to be $I_{\text{cell}}\sim 10^{9}$ bits. This yields a minimum required information accumulation rate of
[…]
corresponding to an accumulation rate of only about 2 bits per year.
Compared to $R(D)\sim 100$ bits/s (modulo the efficiency factor $\eta$), this is minuscule—but reflects the vast timescale over which life could have emerged. To illustrate: a DNA molecule grows by 2 bits per base pair. At a rate of 100 bits/s, this corresponds to 50 base pairs per second, or more than the considerable $500$ m length of a DNA molecule in a year. The remarkably low threshold for information accumulation at $6.34\times 10^{-8}$ bits/s suggests that abiotic emergence of life should be feasible, even with extremely low efficiency ($\eta\sim 10^{-8}$) or short molecular lifetimes ($D\sim 1$ s).
Drift versus random diffusive process for self-assembly
The minimum information accumulation rate required to assemble a protocell of complexity $I_{\text{protocell}}=10^{9}$ bits within $T_{\text{available}}=500$ million years is $v_{\text{min}}=R_{\text{min}}=6.34\times 10^{-8}$ bits/s, highlighting that even extremely inefficient or sporadic processes could, in principle, accumulate the required information over geological timescales. However, this estimate assumes a unidirectional, steadily progressive (ballistic or forward-biased) process. What if the information-gathering process behaves more like a random walk—with fluctuations and reversals—rather than directed drift?
Let us consider a stochastic accumulation model: a biased random walk with velocity $v$ and persistence time $\tau$, analogous to “run-and-tumble” dynamics from bacterial chemotaxis [53]. The effective diffusion constant is given by $D_{\text{eff}}=v^{2}\tau/d$, where $d$ is the dimensionality of the search space. If $v=v_{\text{min}}$, then the mean-square displacement (MSD) after time $T_{\text{available}}$ is
$\langle I^{2}_{\text{protocell}}\rangle=(v_{\text{min}}T_{\text{available}})^{2}=2dD_{\text{eff}}T_{\text{available}}=2v_{\text{min}}^{2}\tau T_{\text{available}},$
independent of $d$. Solving for the crossover time $\tau_{c}$ at which this diffusive process accumulates the same information as directed drift yields $\tau_{c}=T_{\text{available}}/2$. Thus, if $\tau<\tau_{c}=250$ Myr, the randomization is too strong, and the process fails to reach the target within $T_{\text{available}}$ (Fig. 3). High persistence—i.e., strong directional memory—is essential.
To illustrate the timescale sensitivity: With $\tau=1$ s, the wait time for assembling a cell becomes $10^{24}$ yr — roughly a hundred trillion universes stacked end to end. Even if $\tau=1$ yr, the required time is still $T^{\prime}\approx 10^{17}$ yr, about ten million times the universe’s current age. In other words, without immense persistence, life’s emergence becomes cosmologically implausible, potentially pointing to alternative mechanisms.
In summary, the plausibility of information accumulation via biased random walks hinges critically on both the persistence time $\tau$ and the effective information accumulation rate $v$. Highly biased or driven processes dramatically outperform fully diffusive ones in crossing the informational threshold required for life.
Figure 3: How driven does a random walk need to be in information space? Time to reach protocell complexity $I_{\text{protocell}}$ as a function of the persistence time $\tau$ for a biased random walk using minimal information accumulation speed $R_{\text{min}}$, given by $T^{\prime}=\langle I_{\text{protocell}}\rangle^{2}/(2v_{\text{min}}^{2}\,\tau)$ (blue line). The available time of $500$ Myr (red dashed line) is reached for a persistence time of $250$ Myr. Also shown is the age of the universe (yellow line).
Material uptake and energy consumption are not limiting
[…]
Using the minimum required information rate $v_{\text{min}}\approx 6.34\times 10^{-8}$ bits/s and again assuming 10 bits per useful molecule, the corresponding capture rate is $v_{\text{min}}/10\approx 6.34\times 10^{-9}$ molecules/s. This implies that, on average, fewer than one molecule needs to be captured every 5 years. Compared to the Berg and Purcell limit of $10^{9}$ molecules/s, this rate is 13 orders of magnitude lower. Hence, even extremely sparse and selective molecular capture would suffice to meet the informational demands of life’s emergence, given enough time.
If a protocell requires $I_{\text{protocell}}\sim 10^{9}$ bits, and each molecule contributes approximately 10 bits, then the total number of molecules needed is $N_{\text{protocell}}=I_{\text{protocell}}/10=10^{8}$ molecules, consistent with estimates for E. coli [55]. Assuming diffusion-limited uptake, the time required to accumulate $N_{\text{protocell}}$ molecules is $\Delta t=N_{\text{protocell}}/J=10^{8}/(2.27\times 10^{9})\approx 0.044\ \text{s}\approx 1.4\times 10^{-9}$ yr. In principle, a protocell could acquire all necessary components in under 50 milliseconds—if every incoming molecule were retained and useful. In reality, however, most molecules would be non-functional, degraded, or unincorporated, making molecular selection and retention—not flux—the true bottlenecks in prebiotic information accumulation.
To estimate the energy required to assemble a system with $I_{\text{protocell}}\sim 10^{9}$ bits, we compare three frameworks: the Landauer limit, biological ATP-driven synthesis, and technological processes such as PCR. First, the Landauer principle states that erasing one bit of information incurs a minimum energy cost of $E_{\text{min}}=k_{B}T\ln 2\approx 2.8\times 10^{-21}$ J/bit at $T\approx 300$ K. Thus, the theoretical minimum energy for assembling the protocell is $E_{\text{Landauer}}\approx 10^{9}\times 2.8\times 10^{-21}=2.8\times 10^{-12}\ \text{J}$. Second, biological energy costs based on ATP hydrolysis provide about $10^{-19}$ J per molecule. If synthesizing one informational molecule (containing $\sim$10 bits) requires one ATP (a conservative assumption), then $E_{\text{ATP}}=10^{8}\times 10^{-19}=10^{-11}\ \text{J}$. Third, technological synthesis via polymerase chain reaction (PCR) consumes about 1–10 J to amplify microgram-scale DNA. A 109-bit construct corresponds to $\sim$50 $\mu$g of DNA, implying $E_{\text{PCR}}\sim 1\ \text{J}$, or roughly $10^{-9}$ J/bit. Even 1 J is a small amount of energy—equivalent to lifting a 100 g apple by 1 meter or powering a smartphone for 1 second.
Hence, minimal life appears to operate within one to two orders of magnitude of the thermodynamic limit, while modern technology (e.g. PCR) is up to $10^{9}$ times less efficient. This contrast highlights the remarkable energetic efficiency of biological systems and provides a thermodynamic baseline for evaluating the feasibility of prebiotic assembly.
Can information accumulate suddenly?
Above, we estimated the information content of a minimal protocell using Kolmogorov complexity, based on detailed quasi-realistic descriptions of cellular processes encoded in modern computational models. This supports a picture of life emerging through a gradual, steady increase in the organization and complexity of matter—from nonliving to living. This is like putting together a gigantic jigsaw puzzle, and appears like the culmination of a reductionistic belief system. And the required persistence times are extreme, even mind boggling.
An alternative possibility is that life arose via a sudden explosive-like transition — by some yet-unknown self-organization principle of collectivessness triumpying over reductionism, that bridged the gap between random chemistry and selective evolution, potentially circumventing the information bottleneck of early self-replication. This idea is not new. It is most well developed in the study of collectively autocatalytic sets, first introduced by Kauffman (1986) [56]. In such networks, each molecule is catalyzed by another member of the set, enabling system-wide replication without requiring any single molecule to self-replicate or the presence of a membrane-enclosed protocell. This framework offers a resolution to Eigen’s paradox: long, information-rich molecules require error correction to persist, but error-correcting machinery itself demands such complexity [57]. Autocatalytic sets enable increasing functional complexity before templated replication emerges. Mathematical models show that as chemical diversity increases, the probability of forming an autocatalytic set undergoes a sharp phase transition—a sudden onset of self-sustaining autocatalysis [58, 59]. This is analogous to a percolation transition in random graphs.
Such transitions may also explain later leaps in complexity, including the emergence of the brain. Put simply, a brain functions as an internal model of the external world, trained by experience and used to predict future outcomes [60] [^1] This predictive capacity has been formalized in the Bayesian brain hypothesis [61], where internal priors are continuously updated in light of new data, approximating posterior beliefs that guide perception and action.. For example, one can model parameter adaptation in a toy brain using the inverse temperature $\beta$ in an Ising-like system, rather than connectivity or size, as the critical control parameter. In such models, evolutionary fitness peaks near a critical value of $\beta$ when environments are dynamic, heterogeneous, and unpredictable [60]. Though phase transitions technically apply to infinite systems, the benefits of sharp transitions often hold in large finite systems as well [60, 62, 63].
In the spirit of this article, we now return to AI. Suppose that chemical reaction networks in a cell can be viewed as a kind of neural network. Then we may ask: how much information must be encoded in such a network to achieve life-like complexity? Recent work on recurrent neural chemical reaction networks (RNCRNs) has shown that artificial recurrent neural networks can be mapped onto chemical reaction networks [64]. These networks are capable of approximating arbitrary dynamical systems governed by ordinary differential equations. Unlike feedforward neural networks, which approximate static input–output mappings, RNCRNs model time-dependent behaviors such as switching in multi-stable systems or oscillations frequently observed in biology.
This suggests that sufficiently complex chemical reaction networks can mimic neural networks—and hence approximate nearly any behavior, from fast signaling to slow gene regulation and cell-cycle control. This relies on the universal approximation theorem, which guarantees that a neural network with sufficient size can approximate any continuous function to arbitrary accuracy. In this view, a collection of interacting molecules becomes a cell once it can approximate adaptive functions required for survival.
Currently, such RNCRNs remain small ($<10$ species or nodes) [64], and it remains unclear how large a network must be to compute the full repertoire of cellular behavior. Still, we can try and estimate the required network size. The universal approximation theorem states that for any continuous real-valued function and any target error $\varepsilon>0$, there exists a neural network that approximates $f$ within $\varepsilon$ on a compact domain. Recent results give an upper bound on the number of neurons $M$ needed to achieve this approximation:
$M(\varepsilon)\leq C\cdot\varepsilon^{-n/k},$ (4)
where $n$ is the input dimension, $k$ is the smoothness of the function class, and $C$ is a constant depending on $f$. Smoother functions (higher $k$) require fewer neurons. However, as the input dimension $n$ increases, approximation becomes harder—the well-known curse of dimensionality.
How does this network size map to information content? If each parameter (e.g., weight or bias) must be specified with $p$ bits of precision, then the total information required to define the network is
$I_{\text{NN}}\approx M\cdot p.$ (5)
Combining Eqs. 4 and 5, we obtain:
$I_{\text{protocell}}(\varepsilon)\sim p\cdot\varepsilon^{-n/k}.$
This links approximation error $\varepsilon$ directly to informational complexity, assuming uniform precision $p$. Note that this estimate applies only to the dynamic regulatory core of the protocell—not its full genomic, structural, or metabolic description. This framework offers a principled way to estimate the minimal informational content of life-like systems, constrained by approximation theory and encoding limits. When combined with rate-distortion theory, it may help formalize how much structure, memory, and robustness a protocell must have to be functionally complete. However, there is no prediction of a sudden transition to higher complexity.
Let us now take a different route and investigate how large and connected a network must be for computation to emerge spontaneously. A linear threshold model exhibits a percolation transition at a critical mean degree $z_{c}=1$. This means that when, on average, each node is linked to one other, a giant connected component (GCC) for nontrivial computation appears [65]. A minimal network of approximately $10^{4}$ components—if sufficiently connected and properly balanced between activation and inhibition—can begin to compute complex Boolean functions. This supports the idea that a protocell capable of decision-making may require only $10^{4}$ interacting components. If each contributes 10 bits, the total is just $10^{5}$$10^{6}$ bits—well below the earlier estimate of $10^{9}$ bits. But here, we consider only logical capacity, not structural or biochemical implementation, making this a lower bound.
Taken together, sufficiently complex chemical reaction networks can mimic the dynamics of recurrent neural networks, or make the emergence of complex logic functions inevitable. This raises the intriguing possibility that biological systems—even at the molecular scale—approach universal computation in their expressive power. Indeed, chemical kinetics, when properly structured, can implement universal computation [66]. More concretely, chemical reaction networks governed by mass-action kinetics can simulate arbitrary computations and are thus Turing complete [67].
Discussion
Setting aside the statistical fluke argument in an infinite universe, we have explored the feasibility of protocell self-assembly on early Earth. A minimal protocell of complexity $I_{\text{protocell}}\sim 10^{9}$ bits could, in principle, emerge abiotically within Earth’s available timespan ($\sim 500$ Myr)—but only if a tiny fraction of prebiotic interactions ($\eta\sim 10^{-8}$) are persistently retained over vast stretches of time. We approached this question through a conceptual framework combining rate-distortion theory, chemical entropy estimates, and Kolmogorov complexity derived from whole-cell and AI models.
To better understand what must be true for abiotic life to succeed, we proposed a “requirements checklist”: (i) some degree of physical or chemical bias (e.g., compartments, cycles, autocatalytic networks), (ii) sufficient persistence time if information accumulates via a biased random walk through chemical space, and (iii) protection and reuse of functional molecules. If steady progression is too slow, a phase transition may have enabled a more abrupt leap in complexity. Either way, a purely random soup is too lossy—some form of prebiotic informational structure must precede Darwinian evolution.
Naturally, the framework has limitations. Our estimates—of entropy, distortion time, and protocell complexity—are crude at best, drawn from analogies rather than direct measurement. The entropy of the prebiotic chemical pool ($H_{\text{prebiotic}}$) is inferred from analogues like meteorites and Titan’s atmosphere, and the estimated information content of a minimal cell ($I_{\text{protocell}}$) is based on compressed models rather than any physical lower bound. Still, the exercise offers a quantitative scaffold for framing the problem—and perhaps guiding future investigations.
Beyond parameter uncertainty, the deeper challenge is mechanistic: even if the information rate is feasible, the route remains opaque. Where did the directionality (drift $v$ or persistence $\tau$) come from? What structures or environmental constraints enabled long-term memory or error suppression without evolved proofreading? The puzzle deepens as timelines shift: from 3.465 Gy microfossils in Western Australia [15, 16, 17] to a LUCA possibly living around 4.2 Gy ago [18], close to the formation of liquid water at 4.404 Gy [13]. But LUCA is just a branching point. Substantial evolutionary development—possibly involving other, now-extinct lineages—likely occurred between the first protocell and LUCA.
Which returns us, cautiously but irresistibly, to the question: Was Earth terraformed, or did order coalesce from chaos under the silent governance of physics? Today, humans seriously contemplate terraforming Mars or Venus in scientific journals [68, 69]. If advanced civilizations exist, it is not implausible they might attempt similar interventions—out of curiosity, necessity, or design. Still, Occam’s razor weighs in: abiotic evolution, however slow and strange, remains a viable (if mind-bending) explanation. Invoking terraforming adds explanatory complexity without constraint. And while we cannot prove that abiogenesis is inevitable, it remains consistent with thermodynamics.
If life is the ultimate emergent phenomenon, perhaps it resists prediction not because the physics is wrong, but because the framework is incomplete. As Laughlin has argued, emergence defies reductionism not in defiance of physics, but in fulfillment of it [70, 71]. The reaction path may be immensely long, contingent, and distributed across scales—too delicate to reconstruct after four billion years. In such a landscape, quantitative feasibility may be the most we can achieve. But feasibility, even without full predictability, can still inform what is physically possible—and that, for now, is enough to keep the question alive.
Looking ahead, we advocate for more empirical, exploratory work in the spirit of Miller and Urey—to better constrain $\tau$, $D$, and prebiotic entropy. We suggest focusing on mechanisms of memory and information retention in prebiotic systems. For instance, in complex and stochastic networks, do certain configurations converge to robust attractor dynamics in molecular or ecological state space? Might this make evolution more deterministic than traditionally thought, as hinted at by recent outside-of-the-box experiments [72, 73]?
AI may itself play a decisive role. Beyond modeling chemical systems, it may help reverse-engineer candidate pathways, identify attractor landscapes, or uncover deep statistical regularities that escape human intuition. If life is a form of physical computation, as some have proposed [74], then AI might soon help identify how natural chemistry can become computationally powerful enough to self-organize. (Personally, I am skeptical that a generative AI can move from predicting the next word to proving the origin of life, bootstrapping us out of our existential dilemma.) Whether or not AI ultimately solves the origin-of-life problem, it may well turn out to be the next Cohen’s microscope (or telescope), just better [75]. Certainly, it has begun to rewrite the language in which biological complexity can be described.
We end with a note of caution. There is a real possibility that, in seeking to understand life’s origin, we become a living parable of Gödel’s incompleteness and Turing’s undecidability—systems entangled in their own logic, unable to fully explain themselves. Perhaps life cannot prove its own existence; perhaps the AI we task with the job will run forever, or halt with an answer as cryptic as “42.” To avoid standing dumbstruck—like an ape before a lightning-struck fire—we must ensure that our tools, however powerful, can still speak in terms we understand. Otherwise, we risk becoming spectators of intelligence we cannot follow, let alone guide.
“The miracle of the appropriateness of the language of mathematics for the formulation of the laws of physics is a wonderful gift which we neither understand nor deserve.” — Eugene P. Wigner [76]
Acknowledgements
We thank the Physics of Life Network of Excellence at Imperial College, for its stimulating seminars, journal club, and workshops. The latter included ”Physics of emergent behaviour III: from origin of life to multicellularity” in 2021, among others. We also acknowledge the lack of funding, which, while generally detrimental, created space for pondering the fundamental questions of this article.